OpenStudy (anonymous):

–2x + 1 ≤ –11 please explain i got the first one i posted but this one looks even more confuseing :\

hero (hero):

What would you do if -2x + 1 = -11 ?

OpenStudy (anonymous):

As Hero said, just solve it in exactly the same way: group like terms and simplify.

OpenStudy (anonymous):

ummmm added -2 and 1 and got -1 then subtracted -11 from -1 and got 10

hero (hero):

–2x + 1 ≤ –11 -2x ≤ - 11-1 -2x ≤ - 12 12 ≤ 2x 12/2 ≤ x 6 ≤ x

OpenStudy (anonymous):

It is still a litle confusing :\ sorry u dont have to help me if i am to much of a trouble........

hero (hero):

I'm sorry to hear that.

OpenStudy (anonymous):

It is ok u tried atleast

hero (hero):

Maybe this might help... When adding two negative numbers do this: -11 - 1 = -(11 + 1) = -(12) = -12

OpenStudy (anonymous):

Oh ok thanks helped a lil bit :)

hero (hero):

yw